OpenStudy (anonymous):

what is the area of a triangle with a base of 19 inches and a hieght of 11 inches

OpenStudy (anonymous):

So if the formula is A = 1/2bh then A = Area B = base H = height So: A = we don't know yet B = 19 inches H = 11 inches So: A = (1/2) * 19 * 11

OpenStudy (anonymous):

A = (1/2) * 19 * 11 A = (1/2) * 209 A = 104.5

OpenStudy (anonymous):

@joprdan do you get it?